# Julian Martinez-Iriarte, and Gabriel Montes-Rojas

### v3

@@ 4.2 Asymptotic properties
$\Sigma_{\phi}:=\begin{bmatrix}E[Z_{i}Z_{i}^{\prime}]&E[Z_{i}X_{i}^{\prime}]&0\\ E[X_{i}Z_{i}^{\prime}]&E[X_{i}X_{i}^{\prime}]&0\\ 0&0&2E[\nabla_{\gamma}h(X_{i},Z_{i};\gamma)\nabla_{\gamma}h(X_{i},Z_{i};\gamma)^{\prime}h(X_{i},Z_{i};\gamma)^{2}]\\ \end{bmatrix},$
and
$M_{i}=\begin{bmatrix}Z_{i}\left(D_{i}-Z_{i}^{\prime}\pi_{1}-X_{i}^{\prime}\pi_{2}\right)\\ X_{i}\left(D_{i}-Z_{i}^{\prime}\pi_{1}-X_{i}^{\prime}\pi_{2}\right)\\ \nabla_{\gamma}h(X_{i},Z_{i};\gamma)\left(V_{i}^{2}-1\right)h(X_{i},Z_{i};\gamma)^{3}\end{bmatrix}.$
The next theorem contains the limiting distribution of the proposed two-step CF estimator of $\alpha$. This provides the details for the asymptotics for estimator of interest, which could be interpreted as an application of OLS with generated regressors. We consider the following assumption:
Assumption 8.
Let $W(\phi)$ be all the regressors in (11) (except $D$), and define $R_{i}(\phi):=(D_{i},W_{i}(\phi)^{\prime})^{\prime}$. Assume that uniformly in $\phi$ (i) $E[R_{i}(\phi)R_{i}^{\prime}(\phi)]$ is non-singular; (ii) a (uniform) law of large numbers holds for $\frac{1}{n}\sum_{i=1}^{n}R_{i}(\phi)R_{i}^{\prime}(\phi)$, $\frac{1}{n}\sum_{i=1}^{n}(U_{i}J_{\phi,i}-R_{i}(\phi)\alpha^{\prime}J_{\phi,i})$, where $J_{\phi}$ is the Jacobian of $\phi\mapsto R(\phi)$, $J_{\phi}=\nabla_{\phi}R(\phi)$; and (iii) $E[R^{2}U^{2}]<\infty$.
Assumption 8 $(i)$ ensures that the CF additional terms to the structural equation are not perfectly collinear to the main regressors, and Assumptions 8 $(ii)$ and $(iii)$ are required for the asymptotic representations using the influence functions. [^6] Sufficient conditions for a uniform law of large numbers can be found in Lemma 2.4 in Newey and McFadden (1994). Essentially, it depends on a dominating function for $R(\theta)$ or alternatively $h(x,z;\gamma)$.
Theorem 1.
Under Assumptions 1-8, $\widehat{\alpha}$ is asymptotically normal with influence function given by
$\sqrt{n}(\widehat{\alpha}-\alpha)$ $=\Sigma_{\alpha}^{-1}\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\psi_{i}+o_{p}(1),$
where $\Sigma_{\alpha}:=E[RR^{\prime}]$, $\psi_{i}:=R_{i}U_{i}+E[UJ_{\phi}-R(\phi)\alpha^{\prime}J_{\phi}]\Sigma_{\phi}^{-1}M_{i}$.
From the above result, we obtain
$\sqrt{n}(\widehat{\alpha}_{1}-\alpha_{1})\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,\Omega_{1}),$
where $\Omega_{1}$ is the $(1,1)$ element of the limiting variance of $\sqrt{n}(\widehat{\alpha}-\alpha)$, denoted by $\Omega$, and is given by
$\Omega=\Sigma_{\alpha}^{-1}\Psi\Sigma_{\alpha}^{-1}$
with $\Psi=E[\psi\psi^{\prime}]$.
Estimation of $\Omega_{1}$ can be carried out by resorting to the sample counterpart of $\Omega$ as following:
$\widehat{\Omega}$ $=\widehat{\Sigma}_{\alpha}^{-1}\left[\frac{1}{n}\sum_{i=1}^{n}\left(\hat{\psi}_{i}\hat{\psi}_{i}^{\prime}\right)\right]\widehat{\Sigma}_{\alpha}^{-1},$
where
$\widehat{\Sigma}_{\alpha}$ $=\frac{1}{n}\sum_{i=1}^{n}\widehat{R}_{i}\widehat{R}_{i}^{\prime},$
$\widehat{R}_{i}$ $=R_{i}(\widehat{\phi}),$
$\widehat{U}_{i}$ $=Y_{i}-D_{i}\widehat{\alpha}_{1}-W_{i}(\widehat{\phi})^{\prime}\widehat{\alpha}_{w},$
$\hat{\psi}_{i}$ $=\hat{R}_{i}\hat{U}_{i}+\left[\frac{1}{n}\sum_{i=1}^{n}(\hat{U}_{i}\nabla_{\phi}R_{i}(\hat{\phi})-\hat{R}_{i}\hat{\alpha}^{\prime}\nabla_{\phi}R_{i}(\hat{\phi}))\right]\hat{\Sigma}_{\phi}^{-1}\hat{M}_{i},$
@@ 5 Monte Carlo Simulations
$Y$ $=\alpha_{1}D+\alpha_{2}X+g(D,X)\varepsilon,$ (12)
$D$ $=\pi_{1}Z+\pi_{2}X+h(Z,X)V,$ (13)
@@
$g(D,X)$ $=\delta_{1}D+\delta_{2}D^{2}+\delta_{3}X,$ (15)
$h(Z,X)$ $=\gamma_{1}Z+\gamma_{2}X.$ (16)
@@
For the CF models we consider one type of skedastic function that matches the data generating process (16) given in Romano and Wolf (2017), as given in equation (8), $h(X,Z;\gamma)^{2}=|Z^{\prime}|\gamma_{1}+|X^{\prime}|\gamma_{2}=\gamma_{0}+|Z^{\prime}|\gamma_{1}$, such that $\hat{h}=\sqrt{h(X,Z;\hat{\gamma})^{2}}$. Let $\check{V}$ be the residuals of the first-stage regression of $D$ on $Z$ and a constant, then $h^{2}$ is estimated as the regression of $\check{V}^{2}$ on $|Z|$ and a constant.
@@
The results in Tables 1 and 2 clearly indicate that OLS is biased and that 2SLS is also biased in the presence of endogenous heteroskedasticity (i.e. $\delta_{1}\neq 0$ and/or $\delta_{2}\neq 0$). Moreover, note that the parameter $\gamma_{1}$, i.e. heteroskedasticity in the first-stage, aggravates the 2SLS bias. However, the proposed CF estimators provide unbiased results. Note that this depends on the functional form of the heteroskedasticity-inducing function $g$. When $\delta_{2}=0$, both CF1 and CF2 works, but when $\delta_{2}=0.2$, only CF2 completely eliminates the bias. Overall, this indicates that the CF approach eliminates the bias, but nonlinear terms may be needed for different functional forms of the heteroskedasticity inducing functions $g$.
@@
$n$ $\delta_{1}$ $\delta_{2}$ Bias Var. Bias Var. Bias Var. Est.Var Cov.95%
OLS 2SLS CF1 with $\hat{V}+\hat{V}\times D$
250 0 0 0.733 0.004 -0.016 0.026 -0.011 0.026 0.024 0.943
500 0 0 0.734 0.002 -0.006 0.011 -0.004 0.011 0.020 0.957
1000 0 0 0.735 0.001 0.000 0.005 0.001 0.005 0.006 0.958
250 0 0.2 1.806 0.060 0.368 0.216 0.382 0.192 0.325 0.828
500 0 0.2 1.805 0.030 0.382 0.104 0.389 0.094 0.093 0.714
1000 0 0.2 1.806 0.015 0.387 0.052 0.388 0.047 0.047 0.556
250 1 0 2.056 0.080 -0.037 0.337 -0.018 0.308 0.300 0.945
500 1 0 2.056 0.038 -0.010 0.151 -0.002 0.141 0.146 0.955
1000 1 0 2.051 0.019 -0.013 0.078 -0.008 0.071 0.072 0.951
250 1 0.2 3.131 0.216 0.372 0.777 0.403 0.693 2.236 0.894
500 1 0.2 3.123 0.105 0.373 0.368 0.384 0.327 0.340 0.888
1000 1 0.2 3.122 0.054 0.384 0.182 0.392 0.162 0.170 0.832
OLS 2SLS CF2 with $\hat{V}+\hat{V}\times D+\hat{V}\times D^{2}$
250 0 0 0.730 0.004 -0.010 0.024 -0.006 0.026 0.032 0.934
500 0 0 0.732 0.002 -0.007 0.011 -0.006 0.012 0.012 0.946
1000 0 0 0.735 0.001 0.002 0.006 0.003 0.006 0.006 0.942
250 0 0.2 1.801 0.058 0.365 0.204 -0.020 0.164 0.172 0.951
500 0 0.2 1.804 0.029 0.385 0.105 -0.008 0.085 0.084 0.952
1000 0 0.2 1.807 0.015 0.391 0.050 -0.003 0.043 0.042 0.950
250 1 0 2.057 0.075 -0.047 0.309 -0.032 0.299 0.336 0.950
500 1 0 2.050 0.039 -0.015 0.164 -0.007 0.160 0.148 0.943
1000 1 0 2.048 0.019 -0.015 0.085 -0.008 0.081 0.074 0.942
250 1 0.2 3.127 0.208 0.344 0.754 -0.024 0.645 0.702 0.946
500 1 0.2 3.109 0.112 0.355 0.396 -0.029 0.334 0.333 0.949
1000 1 0.2 3.119 0.057 0.383 0.188 -0.025 0.166 0.165 0.946
@@
$n$ $\delta_{1}$ $\delta_{2}$ Bias Var. Bias Var. Bias Var. Est.Var Cov.95%
OLS 2SLS CF1 with $\hat{V}+\hat{V}\times D$
250 0 0 0.614 0.003 -0.021 0.025 -0.009 0.023 0.116 0.943
500 0 0 0.612 0.001 -0.011 0.011 -0.005 0.011 0.011 0.953
1000 0 0 0.612 0.001 -0.004 0.006 -0.001 0.005 0.005 0.954
250 0 0.2 1.934 0.090 0.796 0.321 0.622 0.371 0.346 0.748
500 0 0.2 1.932 0.047 0.830 0.159 0.659 0.183 0.179 0.624
1000 0 0.2 1.942 0.025 0.850 0.086 0.678 0.101 0.094 0.379
250 1 0 1.835 0.066 0.305 0.337 -0.038 0.397 1.295 0.946
500 1 0 1.836 0.033 0.328 0.165 -0.017 0.192 0.191 0.945
1000 1 0 1.834 0.017 0.340 0.082 -0.017 0.097 0.094 0.946
250 1 0.2 3.147 0.271 1.135 1.059 0.620 1.185 1.071 0.864
500 1 0.2 3.167 0.149 1.182 0.487 0.663 0.553 0.564 0.838
1000 1 0.2 3.160 0.074 1.221 0.258 0.697 0.297 0.286 0.723
OLS 2SLS CF2 with $\hat{V}+\hat{V}\times D+\hat{V}\times D^{2}$
250 0 0 0.612 0.002 -0.016 0.025 -0.006 0.026 0.053 0.946
500 0 0 0.614 0.001 -0.006 0.012 -0.002 0.012 0.012 0.944
1000 0 0 0.613 0.001 -0.006 0.006 -0.002 0.006 0.006 0.949
250 0 0.2 1.928 0.098 0.803 0.344 -0.029 0.289 0.649 0.940
500 0 0.2 1.940 0.052 0.841 0.166 -0.018 0.140 0.136 0.946
1000 0 0.2 1.940 0.024 0.845 0.083 -0.011 0.075 0.069 0.937
250 1 0 1.823 0.070 0.298 0.348 -0.044 0.406 2.308 0.939
500 1 0 1.830 0.033 0.322 0.165 -0.027 0.196 0.192 0.950
1000 1 0 1.834 0.017 0.332 0.079 -0.023 0.093 0.095 0.954
250 1 0.2 3.126 0.302 1.084 1.091 -0.097 1.062 1.057 0.940
500 1 0.2 3.140 0.137 1.138 0.513 -0.060 0.526 0.491 0.942
1000 1 0.2 3.176 0.075 1.221 0.251 0.011 0.252 0.247 0.952
@@
$n$ $\delta_{1}$ $\delta_{2}$ Bias Var. Bias Var. Bias Var. Est.Var Cov.95%
OLS 2SLS CF1 with $\hat{V}+\hat{V}\times D$
250 0 0 0.000 0.003 0.001 0.012 0.001 0.012 0.011 0.945
500 0 0 0.000 0.002 -0.001 0.006 -0.001 0.006 0.006 0.941
1000 0 0 0.000 0.001 -0.001 0.003 -0.001 0.003 0.003 0.953
250 0 0.2 0.002 0.028 -0.001 0.098 -0.002 0.097 0.093 0.931
500 0 0.2 -0.005 0.015 -0.001 0.049 -0.001 0.049 0.046 0.939
1000 0 0.2 -0.003 0.007 -0.003 0.025 -0.003 0.025 0.024 0.942
250 1 0 -0.002 0.036 -0.010 0.152 -0.009 0.151 0.148 0.940
500 1 0 0.003 0.019 0.007 0.078 0.006 0.077 0.075 0.943
1000 1 0 0.000 0.010 0.006 0.040 0.006 0.040 0.037 0.936
250 1 0.2 -0.006 0.103 -0.009 0.379 -0.010 0.376 0.364 0.939
500 1 0.2 -0.004 0.051 -0.010 0.188 -0.011 0.187 0.176 0.944
1000 1 0.2 0.005 0.027 -0.004 0.094 -0.004 0.094 0.090 0.944
OLS 2SLS CF2 with $\hat{V}+\hat{V}\times D+\hat{V}\times D^{2}$
250 0 0 0.001 0.003 0.001 0.012 0.002 0.014 0.013 0.947
500 0 0 0.001 0.001 0.002 0.006 0.002 0.007 0.007 0.943
1000 0 0 0.000 0.001 0.001 0.003 0.001 0.003 0.003 0.952
250 0 0.2 0.007 0.028 -0.007 0.096 -0.013 0.107 0.134 0.928
500 0 0.2 0.000 0.014 0.002 0.047 0.004 0.050 0.049 0.943
1000 0 0.2 0.002 0.007 0.000 0.023 0.000 0.024 0.025 0.952
250 1 0 0.006 0.039 0.024 0.163 0.026 0.179 0.315 0.936
500 1 0 -0.004 0.020 0.006 0.078 0.012 0.086 0.082 0.942
1000 1 0 0.000 0.009 0.003 0.039 0.005 0.043 0.041 0.951
250 1 0.2 -0.003 0.102 -0.012 0.365 -0.007 0.389 0.369 0.937
500 1 0.2 0.000 0.049 -0.005 0.181 -0.014 0.199 0.189 0.939
1000 1 0.2 -0.001 0.026 -0.009 0.091 -0.015 0.099 0.095 0.938
@@
$n$ $\delta_{1}$ $\delta_{2}$ Bias Var. Bias Var. Bias Var. Est.Var Cov.95%
OLS 2SLS CF1 with $\hat{V}+\hat{V}\times D$
250 0 0 0.000 0.002 0.002 0.012 0.002 0.011 0.024 0.956
500 0 0 0.000 0.001 0.001 0.006 0.000 0.006 0.005 0.949
1000 0 0 0.000 0.000 0.000 0.003 0.000 0.003 0.003 0.948
250 0 0.2 -0.002 0.039 -0.006 0.151 -0.006 0.181 0.160 0.923
500 0 0.2 -0.001 0.020 -0.001 0.073 -0.002 0.091 0.086 0.935
1000 0 0.2 0.001 0.011 0.005 0.038 0.006 0.047 0.045 0.941
250 1 0 0.003 0.034 0.008 0.177 0.011 0.215 1.039 0.937
500 1 0 -0.002 0.017 -0.003 0.086 -0.003 0.104 0.099 0.937
1000 1 0 0.001 0.009 0.003 0.043 0.005 0.051 0.050 0.940
250 1 0.2 0.000 0.120 0.004 0.494 0.000 0.603 0.567 0.936
500 1 0.2 -0.006 0.061 -0.013 0.244 -0.016 0.301 0.291 0.944
1000 1 0.2 0.004 0.032 0.006 0.123 0.007 0.154 0.152 0.942
OLS 2SLS CF2 with $\hat{V}+\hat{V}\times D+\hat{V}\times D^{2}$
250 0 0 0.001 0.002 0.005 0.012 0.004 0.013 0.048 0.951
500 0 0 0.000 0.001 0.001 0.006 0.002 0.007 0.006 0.945
1000 0 0 0.000 0.000 -0.001 0.003 0.000 0.003 0.003 0.943
250 0 0.2 -0.003 0.038 0.005 0.139 0.012 0.179 0.161 0.937
500 0 0.2 -0.001 0.020 -0.002 0.072 0.006 0.095 0.086 0.933
1000 0 0.2 0.000 0.009 0.001 0.035 0.007 0.050 0.045 0.927
250 1 0 -0.001 0.034 -0.003 0.173 -0.002 0.228 0.217 0.943
500 1 0 -0.002 0.017 -0.008 0.082 -0.011 0.110 0.108 0.943
1000 1 0 0.001 0.009 0.004 0.040 0.006 0.054 0.054 0.948
250 1 0.2 -0.003 0.132 -0.011 0.513 -0.010 0.637 0.713 0.926
500 1 0.2 -0.004 0.062 -0.006 0.257 -0.004 0.339 0.304 0.931
1000 1 0.2 -0.006 0.031 -0.008 0.125 -0.005 0.162 0.155 0.946

### v4

@@ 4.3 Inference
$\widehat{U}_{i}$ $=Y_{i}-D_{i}\widehat{\alpha}_{1}-W_{i}(\widehat{\phi})^{\prime}\widehat{\alpha}_{w}=Y_{i}-\widehat{R}_{i}^{\prime}\widehat{\alpha}$,
$\hat{\psi}_{\phi,i}$ $=\hat{\Sigma}_{\phi}^{-1}\hat{m}_{i},$
@@
Next, we state a result establishing the consistency of the above estimator.
Lemma 3.
If Assumptions 7 and 8 hold, $E[\|R\|^{4}]<\infty$, $E[\|J_{\phi}\|^{4}]<\infty$, $E[U^{4}]<\infty$, $E[\|Z\|^{4}]<\infty$, $E[\|X\|^{4}]<\infty$$E[|h(X,Z;\gamma)V|^{4}]<\infty$, $c_{i}(\phi):=\nabla_{\gamma}h(X_{i},Z_{i};\gamma)\left(V_{i}^{2}-1\right)h(X_{i},Z_{i};\gamma)^{3}$ is differentiable in $\phi$ with derivative $\nabla_{\phi}c_{i}(\phi)$, $E\big[\|\nabla_{\phi}c_{i}(\phi)\|^{2}\big]<\infty$, and satisfies a uniform law of large numbers, then $\widehat{\Omega}_{\alpha}\overset{p}{\to}\Omega_{\alpha}.$
Given the result in Lemma 3 and Theorem 1, inference is standard. A $(1-\theta)\%$ confidence interval for $\alpha_{1}$ can be constructed as $\widehat{\alpha}_{1}\pm Z_{\theta/2}\sqrt{\widehat{\Omega}_{\alpha,(1,1)}}$, where $\widehat{\Omega}_{\alpha,(1,1)}$ is the $(1,1)$ element of $\widehat{\Omega}_{\alpha}$, and $Z_{\theta/2}$ is the corresponding critical value ($\theta/2$ area to the left) from the standard normal distribution.
Moreover, given the asymptotic normality in Theorem 1, general hypotheses on the parameter $\alpha_{1}$, other more generally $\alpha$, can be accommodated by simple Wald-type tests. The Wald process and associated limiting theory provide a natural foundation for the hypothesis $R\alpha=r$ when $r$ is known, with a Chi-square limiting distribution. Non-linear testing can be easily performed by employing the Delta method.
@@ 5 Monte Carlo Simulations
$Y$ $=D\alpha_{1}+X\alpha_{2}+g(D,X)\varepsilon,$ (11)
$D$ $=Z\pi_{1}+X\pi_{2}+h(Z,X)V,$ (12)
@@
$g(D,X)$ $=D\delta_{1}+D^{2}\delta_{2}+X\delta_{3},$ (14)
$h(Z,X)$ $=Z\gamma_{1}+X\gamma_{2}.$ (15)
@@
[…]
@@
The results in Tables 1 and 2 clearly indicate that OLS is biased and that 2SLS is also biased in the presence of endogenous heteroskedasticity (i.e. $\delta_{1}\neq 0$ and/or $\delta_{2}\neq 0$). Moreover, note that the parameter $\gamma_{1}$, i.e. heteroskedasticity in the first-stage, aggravates the 2SLS bias. However, the proposed CF estimators provide unbiased results. Note that this depends on the functional form of the heteroskedasticity-inducing function $g$, but not on correctly specifying $h$. When $\delta_{2}=0$, both CF1 and CF2 works, but when $\delta_{2}=0.2$, only CF2 completely eliminates the bias. Overall, this indicates that the CF approach eliminates the bias, but nonlinear terms may be needed for different functional forms of the heteroskedasticity inducing functions $g$.
@@
$n$ $\delta_{1}$ $\delta_{2}$ OLS 2SLS CF1: $\hat{V}+\hat{V}\times D$ CF2: $\hat{V}+\hat{V}\times D+\hat{V}\times D^{2}$
Bias Var. Bias Var. Bias Var. Est.Var Cov.95% Bias Var. Est.Var Cov.95%
250 0 0 0.733 0.004 -0.016 0.026 -0.011 0.026 0.024 0.943 -0.006 0.026 0.032 0.934
500 0 0 0.734 0.002 -0.006 0.011 -0.004 0.011 0.020 0.957 -0.006 0.012 0.012 0.946
1000 0 0 0.735 0.001 0.000 0.005 0.001 0.005 0.006 0.958 0.003 0.006 0.006 0.942
250 0 0.2 1.806 0.060 0.368 0.216 0.382 0.192 0.325 0.828 -0.020 0.164 0.172 0.951
500 0 0.2 1.805 0.030 0.382 0.104 0.389 0.094 0.093 0.714 -0.008 0.085 0.084 0.952
1000 0 0.2 1.806 0.015 0.387 0.052 0.388 0.047 0.047 0.556 -0.003 0.043 0.042 0.950
250 1 0 2.056 0.080 -0.037 0.337 -0.018 0.308 0.300 0.945 -0.032 0.299 0.336 0.950
500 1 0 2.056 0.038 -0.010 0.151 -0.002 0.141 0.146 0.955 -0.007 0.160 0.148 0.943
1000 1 0 2.051 0.019 -0.013 0.078 -0.008 0.071 0.072 0.951 -0.008 0.081 0.074 0.942
250 1 0.2 3.131 0.216 0.372 0.777 0.403 0.693 2.236 0.894 -0.024 0.645 0.702 0.946
500 1 0.2 3.123 0.105 0.373 0.368 0.384 0.327 0.340 0.888 -0.029 0.334 0.333 0.949
1000 1 0.2 3.122 0.054 0.384 0.182 0.392 0.162 0.170 0.832 -0.025 0.166 0.165 0.946
@@
$n$ $\delta_{1}$ $\delta_{2}$ OLS 2SLS CF1: $\hat{V}+\hat{V}\times D$ CF2: $\hat{V}+\hat{V}\times D+\hat{V}\times D^{2}$
Bias Var. Bias Var. Bias Var. Est.Var Cov.95% Bias Var. Est.Var Cov.95%
250 0 0 0.614 0.003 -0.021 0.025 -0.009 0.023 0.116 0.943 -0.006 0.026 0.053 0.946
500 0 0 0.612 0.001 -0.011 0.011 -0.005 0.011 0.011 0.953 -0.002 0.012 0.012 0.944
1000 0 0 0.612 0.001 -0.004 0.006 -0.001 0.005 0.005 0.954 -0.002 0.006 0.006 0.949
250 0 0.2 1.934 0.090 0.796 0.321 0.622 0.371 0.346 0.748 -0.029 0.289 0.649 0.940
500 0 0.2 1.932 0.047 0.830 0.159 0.659 0.183 0.179 0.624 -0.018 0.140 0.136 0.946
1000 0 0.2 1.942 0.025 0.850 0.086 0.678 0.101 0.094 0.379 -0.011 0.075 0.069 0.937
250 1 0 1.835 0.066 0.305 0.337 -0.038 0.397 1.295 0.946 -0.044 0.406 2.308 0.939
500 1 0 1.836 0.033 0.328 0.165 -0.017 0.192 0.191 0.945 -0.027 0.196 0.192 0.950
1000 1 0 1.834 0.017 0.340 0.082 -0.017 0.097 0.094 0.946 -0.023 0.093 0.095 0.954
250 1 0.2 3.147 0.271 1.135 1.059 0.620 1.185 1.071 0.864 -0.097 1.062 1.057 0.940
500 1 0.2 3.167 0.149 1.182 0.487 0.663 0.553 0.564 0.838 -0.060 0.526 0.491 0.942
1000 1 0.2 3.160 0.074 1.221 0.258 0.697 0.297 0.286 0.723 0.011 0.252 0.247 0.952
@@
$n$ $\delta_{1}$ $\delta_{2}$ OLS 2SLS CF1: $\hat{V}+\hat{V}\times D$ CF2: $\hat{V}+\hat{V}\times D+\hat{V}\times D^{2}$
Bias Var. Bias Var. Bias Var. Est.Var Cov.95% Bias Var. Est.Var Cov.95%
250 0 0 0.000 0.003 0.001 0.012 0.001 0.012 0.011 0.945 0.002 0.014 0.013 0.947
500 0 0 0.000 0.002 -0.001 0.006 -0.001 0.006 0.006 0.941 0.002 0.007 0.007 0.943
1000 0 0 0.000 0.001 -0.001 0.003 -0.001 0.003 0.003 0.953 0.001 0.003 0.003 0.952
250 0 0.2 0.002 0.028 -0.001 0.098 -0.002 0.097 0.093 0.931 -0.013 0.107 0.134 0.928
500 0 0.2 -0.005 0.015 -0.001 0.049 -0.001 0.049 0.046 0.939 0.004 0.050 0.049 0.943
1000 0 0.2 -0.003 0.007 -0.003 0.025 -0.003 0.025 0.024 0.942 0.000 0.024 0.025 0.952
250 1 0 -0.002 0.036 -0.010 0.152 -0.009 0.151 0.148 0.940 0.026 0.179 0.315 0.936
500 1 0 0.003 0.019 0.007 0.078 0.006 0.077 0.075 0.943 0.012 0.086 0.082 0.942
1000 1 0 0.000 0.010 0.006 0.040 0.006 0.040 0.037 0.936 0.005 0.043 0.041 0.951
250 1 0.2 -0.006 0.103 -0.009 0.379 -0.010 0.376 0.364 0.939 -0.007 0.389 0.369 0.937
500 1 0.2 -0.004 0.051 -0.010 0.188 -0.011 0.187 0.176 0.944 -0.014 0.199 0.189 0.939
1000 1 0.2 0.005 0.027 -0.004 0.094 -0.004 0.094 0.090 0.944 -0.015 0.099 0.095 0.938
@@
$n$ $\delta_{1}$ $\delta_{2}$ OLS 2SLS CF1: $\hat{V}+\hat{V}\times D$ CF2: $\hat{V}+\hat{V}\times D+\hat{V}\times D^{2}$
Bias Var. Bias Var. Bias Var. Est.Var Cov.95% Bias Var. Est.Var Cov.95%
250 0 0 0.000 0.002 0.002 0.012 0.002 0.011 0.024 0.956 0.004 0.013 0.048 0.951
500 0 0 0.000 0.001 0.001 0.006 0.000 0.006 0.005 0.949 0.002 0.007 0.006 0.945
1000 0 0 0.000 0.000 0.000 0.003 0.000 0.003 0.003 0.948 0.000 0.003 0.003 0.943
250 0 0.2 -0.002 0.039 -0.006 0.151 -0.006 0.181 0.160 0.923 0.012 0.179 0.161 0.937
500 0 0.2 -0.001 0.020 -0.001 0.073 -0.002 0.091 0.086 0.935 0.006 0.095 0.086 0.933
1000 0 0.2 0.001 0.011 0.005 0.038 0.006 0.047 0.045 0.941 0.007 0.050 0.045 0.927
250 1 0 0.003 0.034 0.008 0.177 0.011 0.215 1.039 0.937 -0.002 0.228 0.217 0.943
500 1 0 -0.002 0.017 -0.003 0.086 -0.003 0.104 0.099 0.937 -0.011 0.110 0.108 0.943
1000 1 0 0.001 0.009 0.003 0.043 0.005 0.051 0.050 0.940 0.006 0.054 0.054 0.948
250 1 0.2 0.000 0.120 0.004 0.494 0.000 0.603 0.567 0.936 -0.010 0.637 0.713 0.926
500 1 0.2 -0.006 0.061 -0.013 0.244 -0.016 0.301 0.291 0.944 -0.004 0.339 0.304 0.931
1000 1 0.2 0.004 0.032 0.006 0.123 0.007 0.154 0.152 0.942 -0.005 0.162 0.155 0.946